OpenStudy (anonymous):

You have 5 friends who want to go to a concert with you, but you only have 3 extra tickets. How many possible ways are there to choose 3 of your 5 friends? A. 15 B. 10 C. 60 D. 125

OpenStudy (chriss):

\[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\] or 10

OpenStudy (anonymous):

which one could it be hmm. . .

OpenStudy (chriss):

They are the same. 5 choose 3 equals 10

OpenStudy (anonymous):

ok thank you

OpenStudy (chriss):

\[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\] or as it's read, 5 choose 3 is the method you would use to see how many ways you can choose 3 things out of a total of 5 things.